OpenStudy (anonymous):

Find the probability that a normal variable takes values more than 4/5 standard deviations away from its mean.

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

value = mean + (4/5) SD z = [(mean + (4/5) SD - mean]/SD = 0.800 find cumulative F(z=0.8), then greater than that is P = 1-F(z=0.8). Try it and so will I.

OpenStudy (anonymous):

i got .2119

OpenStudy (anonymous):

am i supposed to multiply it by 2?